Ford has developed a new hybrid car called the Fractal. 31 Fractal cars and 27 competitors cars were road tested to compare miles-per-gallon (mpg) performance. The sample statistics are shown in the table below: Sample 1 Sample 2 (competitior) 27 Sample size Sample mean Sample Standard deviation 31 33.0 mpg 31.0 mpg 20 mpg 2.5 mpg Can we conclude, using a 5% level of significan cer that the miles-per-gallon (mpg) performance of Fractal cars is greater than the miles-per gallon performance of the competitors cars? (Note from the Instructor- Think about this concept for your studying purposes) For you answer in this problem- give the calculated test statistic you would come up with in step 5 of the hypothesis testing. process. Give answer to 3 decimal places Your Answer: Answer
Answer #1

Hello Sir/ Mam

YOUR REQUIRED ANSWER IS TEST STATISTIC = 3.331

H_0 : \mu_1 = \mu_2
H_a : \mu_1 > \mu_2

SE = \sqrt{\frac{2^2}{31} + \frac{2.5^2}{27}} = 0.6004

Test \ Statistic = \frac{\mu_1-\mu_2}{SE}= \frac{33-31}{0.6004}=3.331

I hope this solves your doubt.

Feel free to comment if you still have any query or need something else. I'll help asap.

Do give a thumbs up if you find this helpful.